# Social Media Isn’t Just Instagram: A Youth-Envisioned Platform for Meaningful Social Connections

### v1

@@ 4.2.3. Ambient Co-Presence
4.2.4. Effortful and Intentional Navigation
Physical travel was seen as a key way to make interactions feel more meaningful and intentional. While teleportation options like “spells” (P06) or the “Marauder’s Map” (P21) were seen as the convenient default, several wanted the option of being able to travel around the virtual world, as if they were taking a “nature walk” (P14). As P21 remarked, “Many times… the best memories happen during a walk.”
For some, the value of their redesigned interactions would come from the intentionality they would bring. P02 described engaging in SMH with deliberate purpose rather than habitual use, preferring a slower pace: “Teleporting would be nice, but not in this situation for me… I wouldn’t really want anything to be super fast decisions. I want everything to be very thought through.” Similarly, P05 envisioned navigation in SMH as requiring focused effort:
“[Ideally, SMH would] have some way that you specifically have to search for stuff… where you have to intentionally go in looking for the account or the post versus it just being handed to you… you would have to walk through a town to get there. It’s not challenging to navigate, but you have to intentionally go there to see it. It isn’t something you can do mindlessly.” (P05)
P14 highlighted how such intentional interactions on SMH would feel more “personal” compared to seeing “a random 2D Insta post that they posted 3 months ago.” Participants also noted that SMH’s intentional nature would amplify SMH’s focus on meaningful connections. P05 speculated, “Influencers probably wouldn’t be as keen on something that is more time-consuming… People who are genuinely just trying to connect with others would probably benefit… because it is more personal.”
For others, the fulfillment came from the embodied experience of physical actions. P04 valued “meandering through different people’s spaces,” finding that deliberate exploration heightened their sense of being “more present in somebody’s space.” P09 echoed this, appreciating how moving physically between rooms created a “feeling of time passing.” They explained, “there’s a feeling like you’re actually officially leaving a space as opposed to clicking into a different one.”
4.3. Expressive and Lower-Stakes Sharing
Participants emphasized that SMH should enable rich personal expression, allowing them to authentically and casually reflect their personalities and interests through memory-sharing and customization, from decorating dorm rooms to avatars.
4.3.1. Expression Through Personal Area and Avatar Customization
Participants wanted to create and personalize virtual environments and avatars that reflect their identities, preferences, and aspirations. Unlike MSM that only offers “two modes” (P10)— “light and dark mode[s]” (P10)—P21 envisioned a highly customizable world where users could create their own “mini world” within a larger virtual space, complete with “houses, teleport points, and rooms like a kitchen, a bathroom, a bedroom, or even a school.” They emphasized how this space could be shared with friends, “showing a little part of yourself.” P03 imagined a “dark academia” Hogwarts-inspired room filled with “so many bookshelves and… Victorian architecture,” to reflect their interests. P01 expanded on these ideas, envisioning a house with rooms “dedicated to each of your interests.” They added how that would enable them to showcase the different facets of their identities: “I’m into multiple things—you could have a room for Harry Potter, a room for neuroscience, and a room for another fandom or interest.”
“3D avatar[s]” (P11) and profiles were another key way participants envisioned expressing themselves. P22 highlighted how platforms like Myspace (mys, [n. d.]) allowed for a level of customization that reflected individuality, describing how they enjoyed coding profiles to change “the font and really make it reflect me.” P01 also shared that there is an inherent “joy” in “just that process of creating it, and finding out who you are through that process.” Even just choosing a username was seen as a way to “convey to the world and your friends your own identity… which is a positive feeling usually.” Reflecting on Tumblr, P01 explained how such personalization is “mainly for you,” not just for others to see. P17 felt similarly, stating, “I don’t think it necessarily conveys personal aspects of yourself, but I think it will without you meaning it to.” As they put it, “your brain conjured that up, right? So I think that’s part of you.”
Participants highlighted how expression through choice feels authentic. For example, P01 felt that such “creative expression” could feel “a lot more personal” because “that’s not just like what life has given you… a reflection of your outlook on life.” P13 described this kind of indirect sharing as “mak[ing] social media an extension of everyday interaction,” allowing people to “gain the basic foremost information” about someone without requiring direct conversations. P21 suggested this would make it “easier for people to approach you and say, ‘Hey, I like that, too.”’
P10 explained that being in a 3D, VR environment amplifies this expressive nature, stating that 2D worlds feel more “compressed” and “able to see as much stuff as you can see in 3D.” P16 noted that in VR, others are “basically right in front of you when they’re talking,” making their “mannerisms… clearly easily observable.” They elaborated that one could learn a lot about someone based on “the way they hold themselves… the way they decorate their house… how they choose to dress their in-game character.” This experience feels more authentic, as though meeting with “more real people… compared to the profile pictures.”
In addition to expressiveness, participants appreciated how these features support subtle and low-stakes sharing. P16, who described themselves as “fairly private” and not particularly “public” about their preferences, valued the ability to “share things more subtly.” For instance, they explained that “decorat[ing] a house” or customizing an avatar allows others to “get to know more about my personality… through my style.” They found this form of self-expression to be “more justified” than posting, which felt “embarrassing” and “just doing this for other people.”
4.3.2. Real-Time Sharing: Capturing States and Changes
Another recurring theme was the ability to reflect real-time states through external artifacts in virtual spaces. P04 shared how they might opt into sharing updates on “what I’m doing at the moment. For example, if I’m at the shops, maybe I’d want to share that.” P21 extended this idea to items purchased such as a “new pair of New Balance shoes,” imagining virtual spaces that automatically displayed “recently added or recently stored items.”
Participants envisioned that the blending of physical and virtual selves would allow for a more personalized medium of sharing. P01 imagined their avatar reflecting what outfit they’re wearing in the moment or what song they’re listening to, adding that this could help friends intuitively know “‘Oh. that’s what she’s listening to right now’ or ‘That’s what she’s thinking about.”’ P10 described how frequent updates to virtual rooms could reveal changes in personal interests, providing friends with “an insight into… what I’m thinking.”
Participants also discussed the ability to adapt avatars to reflect real-time emotions or actions. P21 described avatars that “read your emotions… showing [them] on your faces,” while P19 envisioned avatars mimicking real-life gestures. Music was frequently mentioned as a subtle and meaningful way to share current moods.
4.3.3. Sharing Memories: Multisensory and Emotional Experiences
A major theme was the ability to share memories directly with others, like the “Pensieve” (pen, [n. d.]). P04 imagined sharing experiences like “jumping into a pool or taking the first bite of an ice cream sundae” directly, creating more of an emotional connection. Sharing memories could also help reduce misunderstandings by enabling more accurate and nuanced expression, unlike MSM, where communication is mostly “just words.” (P03)
P07 analogically described memory-sharing as “screen sharing, but with your eyes.” P08 highlighted the practicality of memory-sharing, explaining how it could capture transient moments—like crossing a state line—without needing to pull out a camera. P03 highlighted how this would enable them to “just give” memories to their mom in real-time, instead of doing so “after the fact,” where they often found themselves missing the right opportunities.
P04 added that memory-sharing would foster more deliberate consumption of content:
“[SMH] feels more personal and interactive… if I could take more time in a less easy-to-consume format with what others choose to share.… On Instagram, when you click on someone’s page, you see all the posts at once. You don’t go one by one, taking time to read through every comment.… Here, you’d have to take more time to hone in on a memory…… It’s the satisfaction of consciously consuming content from someone you’re close to, being deliberate in viewing what they share, versus just quickly looking and being done.” (P04)
Some participants imagined that these memories would not just be sensory but should also include emotions and perspectives. P02 explained how sharing memories could help others “understand why I said what I said” and prevent “misconstru[al]s.” (P02) For some, memories weren’t just visual; P01 expressed a desire to “transmit smells through social media,” adding another sensory dimension to shared experiences. Given ethical concerns, many felt that only minimal editing, such as “cropping certain parts you don’t want to share” (P01) or versions “sped-up or slowed-down,” (P07) should be allowed.
4.4. Granular, Intuitive, and Fun Privacy Mechanisms
Participants felt that current social media platforms offer limited privacy options, often lacking flexibility and nuance. P16 noted, “It’s kind of a pain to organize, and… right now in Instagram you can put close friends… that’s about it.” They expressed a need for more granular controls, particularly in a platform like SMH, where users might share more authentic aspects of their lives. As P11 explained, “If people are posting more like their genuine real life and their experiences, [privacy] would be a little bit more complicated,” emphasizing the importance of having “flexibility” to control access.
4.4.1. Space-Based Privacy
Participants envisioned using spatial or visual cues to establish privacy boundaries, drawing on intuitions from physical spaces. P03 described public areas like “the downstairs area,” “guest bedroom,” or “game room” as spaces where people could “just come in” without permission, while private spaces like their bedroom required “extra consent” as it was a “safe space.” P13 suggested intermediate zones like a “waiting room” for sharing “semi-personal things” before granting further access.
P13 elaborated on how privacy norms in physical world contexts transfer to the SMH world, stating:
“A bedroom is a respected personal space… If you label it as a bedroom, it just feels more intimate. It’s where you spend most of your time, and it shows your personality the most. It’s not like you can easily change the stuff you have in your room.… If someone were to come over, you can’t just rearrange it in 30 minutes—or 30 seconds—the way you can on Instagram. If you wanted to add someone there, all you’d have to do is remove a couple of things. It’s easier than physically changing a space.” (P13)
While managing visibility across multiple rooms might initially seem more complex, participants believed this approach would feel intuitive and natural. P11 suggested that spatial contexts could simplify privacy controls, stating, “It would make things easier because you’d be able to easily tell who’s allowed and not allowed in certain areas of the house.”
4.4.2. Playful Privacy
Participants envisioned engaging privacy mechanisms imbued with a sense of playfulness and fantasy. P11 proposed a system where houses or specific objects within personal rooms could incorporate elements like a “secret knock” or an object that reveals a “secret passage to… secret rooms” when interacted with in a specific way known only to an exclusive list of friends. Others suggested knowledge-gated access, such as a “secret handshake” (P21) with their friends. P02 drew inspiration from the Marauder’s Map, envisioning a system where users could only see accessible spaces unlocked by summoning secret “spells.” P05 likened playful privacy mechanisms to “a little Easter egg”, stating “There’s that little kid part of everyone” that makes the features enjoyable. They noted that these elements would feel “fun” and “novel” as opposed to MSM’s privacy settings, which often come across as uninspired or confusing, “like a wall of settings where sometimes the title of what the setting is doesn’t even make sense.”
4.4.3. Contextual Privacy
Participants reflected on the challenges of navigating context collapse on MSM (Marwick and Boyd, 2011), and hoped SMH would make contextual disclosure easier:
“On Muggle social media, you’ve got to create a whole new account. But I’m thinking for this app, it just telepathically knows what’s in Muggle area and what’s in Hogwarts area. So you don’t really have to go through the effort of keeping two lives. You can just walk through a portal. And you’re like a new person.” (P02)
P03 proposed a framework where every user has a “base” identity but could customize additional aspects of their representation. P22 imagined a contextual profile where “any group you’re in…you could change your avatar for that,” drawing inspirations Discord where “you can make different profiles for different servers.” P14 highlighted the importance of individual control, suggesting that “each of them should have their own setting, because I feel like everybody has different levels of what they want to keep to themselves.”
Participants also emphasized that contextual privacy goes beyond person-based access control. For example, they suggested varying access based on temporal factors. P14 proposed that “people who don’t follow you… can’t see past versions… I’m cool with you seeing the 6-year-old version of me on this, but I don’t want you to see my 12-year-old version.” The idea of gradual openness was another suggestion:
“if you’re meeting a complete stranger, you wouldn’t want to give them like exactly how you look… But then, after a couple of conversations, you sort of get the gist of how they are, and whether you want them to see you… so then you can turn it on. Or turn it off.” (P21)
4.4.4. Invisibility
Participants were critical of existing platforms for their lack of a robust blocking system. P22 noted that on Instagram, even with contact syncing off, “they might still suggest accounts based on the follow relationships,” exposing a Finsta (Fake + Instagram; secondary account (Huang and Vitak, 2022)) account to people they would rather avoid. They also emphasized the need to block someone thoroughly, including their “alternate accounts,” that “I don’t wanna be able to tell that the person I blocked is even around me.” P14 added that on Instagram, even private accounts allow “a random person” to see personal content such as Story highlights (Bojkov, 2019).
Many users expressed a desire for greater control over their visibility on social media, emphasizing features that would allow them to be effectively “invisible” unless they chose otherwise. P03 envisioned users having “an invisibility cloak” for situations where “you don’t want anyone to see you.” Similarly, P07 proposed a “super private” profile option where others “can’t even search… like you’re not even on the app.”
The ability to prevent unsolicited contact was another recurring theme. P22 suggested a model inspired by Discord, where “unprompted contact” happens only with people “you’ve already interacted with,” and others “can’t even be discovered unless you have something in common.” This approach resonated with P13, who felt strongly that “people should only be able to add you if you meet in person and exchange a QR code.” They reflected on an experience where they were able to find personal information about a waitress online based just on their name, noting “it’s too easy to find people on social media.”
4.4.5. Age-Appropriate Spaces
Participants emphasized the importance of creating age-appropriate spaces for minors. P21, for example, argued “people who are 50” should not be engaging with “people who are 13.” Age verification was widely seen as a critical first step. P16 suggested incorporating “biometrics” into the verification process, while P19 argued that users should “legally prove” their age. Once reliable age verification is in place, participants advocated for creating separate spaces for minors to which older users do not have access. Conversely, P22 proposed that minors should have restrictions, such as being unable to join general spaces like a “town square.”
4.5. Refocusing Interaction Priorities for Meaningful Engagement
Participants expressed a strong desire to refocus their social media experiences on meaningful connections and intentional interactions. This section explores four key themes that emerged from their responses: the need to prioritize genuine connections over algorithmic content recommendations, reducing the influence of celebrities and commercial content, using spatial metaphors to represent relationship strengths, and fostering more intentional and meaningful use of time. These themes reflect participants’ vision for social media platforms that better align with their desires for authentic social connection and purposeful engagement.
4.5.1. Intentional Content Consumption
Participants emphasized the importance of refocusing social media on genuine, personal connections rather than extraneous content. P19 criticized how platforms like Instagram have shifted away from its initial focus:
@@ 4.5.1. Intentional Content Consumption
This sentiment was echoed by P22, who lamented that on Instagram, “half of the timeline is accounts I don’t even follow that were just randomly recommended to me.” They described this experience as “less keeping up with friends and more just getting random stuff shoved in my face.”
On a related note, a major source of frustration was the infinite scrolling design of current platforms. P11 explained, “There’s always more content on social media… this fuels the mindset that… I need to make sure I’m not missing anything.” P14 also criticized the endless content stream on TikTok, which exposes users to “clickbaity” and often harmful content that “doesn’t stimulate [users] at all” and “messes with [users’] attention span.” They expressed a desire to “be more proactive” and “genuinely use [their] brain” instead of “just scrolling mindlessly for an hour.” P09 envisioned a system like an elevator: “You press which floor you want to go to, and then, when you ride that elevator, it just goes straight to that floor… Could [SMH] be sort of like that, where you think about what you want to see before entering that space, and then it shows what you were thinking about?” This sentiment was echoed by several participants who envisioned brain-controlled interfaces that would “automatically show [users] things [they] would really like” (P07) based on their immediate needs or interests.
4.5.2. Reduced Celebrity and Commercial Influence

### v2

@@ 4.2.2. Making Space for Casual Connections Through Low-Pressure, Ambient Social Interaction
4.3. Communal Ambience
SMH was envisioned as a platform designed for intentional, engaging interactions with close friends, making it naturally less prone to issues like social comparison. P09 described it as “the antonym to pressure,” “stress-free,” “community oriented,” “wholesome,” “accessible,” and “accepting.” Unlike MSM, which often feels performative and stressful, participants emphasized features and structures fostering organic, intentional, and personalized interactions, leading to closer friendships. Participants described several design features, such as neighborhood structures, shared gathering spaces, and stronger boundary controls, that could help cultivate this kind of communal, low-pressure environment.
4.3.1. Designing Neighborhood Structures that Reflect Real-World Communities
Participants wanted SMH’s 3D world to mirror the structure and functionality of real-world communities, using familiar layouts like homes, neighborhoods, and shared areas to naturally foster social interaction and relationship-building. P22 described in detail:
“So, in a virtual world, you might have your house, and then you could go to a community board or town square with postings for different groups … From there, you could travel to those group spaces. I don’t know if you’ve ever played or seen Club Penguin, but in that world, you had your igloo to decorate and could visit the town … The town had different buildings for different activities. I’m imagining something like that, where you have your house where people can visit and talk with you, and if you want to meet new people, you can leave your house and go out into the town.” (P22)
Youth envisioned SMH as also reflecting the dynamics of real-world relationships, using physical distances within the SMH landscape to intuitively represent the strength of social ties. Specifically, some participants envisioned SMH landscape to resemble a college campus or a neighborhood, where physical closeness naturally fosters interaction. One participant elaborated:
“And in this magical world, I imagine the people around you in your neighborhood would be from the same house as you, kind of like how people live together in dorms on a college campus. The nearby stores would be the ones your group usually goes to. So you’d share that sense of togetherness because you all go to the same stores or belong to the same house. I think that would help build a real sense of community.” (P03)
Similarly, P01 proposed a tiered layout, where “the top 7 people you’re closest to would be the 7 houses in your cul-de-sac area,” while acquaintances would be positioned farther away, requiring a “little hike” to reach them. This design approach of aligning physical landscapes with emotional closeness was seen as a way to streamline interactions and focus on meaningful relationships. P07 noted how, in short bursts of social media use, they preferred seeing content from their closest friends over “some girl I talked to once three years ago.”
4.3.2. Establishing Third Places for Casual Social Interaction and Serendipitous Connections
Participants imagined that in addition to private homes and personal spaces, SMH neighborhoods would include shared third places where people could casually gather, lowering social barriers and creating opportunities for connection. P05 described these areas as “neutral spaces”, not tied to any specific activity or relationship status, where groups could gather for a range of purposes. They imagined “some places to sit, and some kind of flat table-like surface” that could support casual activities like “doing homework or study, or [playing] some kind of board game.” Such spaces, they explained, would function as “a third, neutral territory” where people could interact more organically, without the friction of managing invitations or switching platforms.
Participants described shared social spaces within neighborhoods, such as common rooms, courtyards, or clubhouses, where residents could casually interact outside their personal spaces. P01 described leaving one’s home to gather in a “courtyard or game room” at the center of the neighborhood. Drawing on examples from Club Penguin  and Poptropica , they imagined this shared space as a “clubhouse… open to anybody,” with themed rooms that cater to different interests, such as a Quidditch field or library, creating opportunities for people to meet others through shared activities. P09 similarly imagined a “virtual dining room” that would allow people to casually talk while doing something familiar, like eating together. Unlike the more segmented house-based accommodation at Hogwarts, this dining space was described as open and inclusive, allowing people to connect across social groups.
Others described third places located in public areas that supported casual, serendipitous interactions. P03 described chance encounters while shopping: “you strike a conversation because you’re both picking up the same shirt… little connections like that happen just because you’re in a public space.” Similarly, P08 suggested a marketplace space where selling items could spark conversations: “‘Hey, look! This is for sale’… have a talk about it… and meet new people.” P21 imagined “small arenas” for “Quidditch games…[where users] can choose to get seated randomly [and] just start talking to the person next to you…and meet new people.” Other suggestions included “public study hall[s]” (P17) where participants could “interact with everybody,” (P17) leading them to experience that, “all of a sudden, you know this person that plays the piano.” (P17)
4.3.3. Prioritizing Close Relationships and Reducing Extraneous Content
To preserve SMH’s focus on meaningful, stress-free interactions, participants emphasized the importance of filtering out irrelevant content, such as celebrity drama or political commentary, and prioritizing updates from close friends and trusted communities. P19 criticized how platforms like Instagram have shifted away from their initial focus: